OpenStudy (anonymous):

Lindsay invested $4500 at 4% interest compounded annually. How much interest will she earn in 10 years? A. $180.00 B. $187.27 C. $1800.00 D. $2161.10

OpenStudy (anonymous):

In order to understand this problem, you have to understand that 4% = 0.04

OpenStudy (anonymous):

If you multiply 4500 by 0.04, you get 180 dollars This is only in 1 year To find the answer in 10 years, you have to multiply 180 by 10 =1800 Does that make sense?

OpenStudy (anonymous):

it might make sense but it is wrong

OpenStudy (anonymous):

to increase a number by \(4\%\) multiply it by \(100\%+4\%=104\%=1.04\) to do it ten times in succession, multiply it by \[(1.04)^{10}\] your job it to compute \[4500\times (1.04)^{10}\] with a calculator

OpenStudy (anonymous):

i get this http://www.wolframalpha.com/input/?i=4500%281.04%29^10

OpenStudy (anonymous):

if you want just the interest, subtract \(4500\) from that number

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

This is compounded interest, not simple interest @satellite73

OpenStudy (anonymous):

right

OpenStudy (anonymous):

the method shown by @KirbyLegs above would be for "simple" interest 4% of the original amount each year the method i wrote is for "compound" interest 4% of the CURRENT amount each year

OpenStudy (anonymous):

That means the answer of ((4500*(1*0.04)= 1st year 1st year total*(1*0.04)= 2nd year And so on and on

OpenStudy (anonymous):

year one \(4500(1.04)\) year two \(4500(1.04)(1.04)=4500(1.04)^2\) year three \(4500(1.04)^3\) year ten \(4500(1.04)^{10}\)

OpenStudy (anonymous):

Yes, but you would take the totals, it would be more work

OpenStudy (anonymous):

i wouldnt care

OpenStudy (anonymous):

Satellite and I both gave answers that are options in the multiple choice problem I'm sure one of us is right. If you are in an algebra class, then I am right If you are in an economics o college level class, Satellite is right Just guessing . . .

OpenStudy (anonymous):

whateve

OpenStudy (anonymous):

The answer was D. I knew the answer, and I could of done it, but to be honest, I was just kinda lazy.